# Inducing Causal World Models in LLMs for Zero-Shot Physical Reasoning

### v3

@@ Abstract
Index Terms:
Large Language Models, World Models, Causal Inference, Physical Reasoning, Zero-Shot Learning, Deep Learning, Artificial Intelligence.
@@ I Introduction
The last few years have witnessed a paradigm shift in artificial intelligence, driven by the advent of Large Language Models (LLMs). Foundational models like BERT [1] and the GPT series [2, 3] have demonstrated remarkable, often emergent, capabilities across a vast spectrum of natural language tasks. Their proficiency is not limited to text; recent multimodal models such as Gemini [4, 5] and Flamingo [6] have extended this power to encompass vision and other modalities, achieving impressive results in areas like visual dialog [7] and video question answering [8]. The capabilities of these models are so extensive that they are often evaluated on large, complex benchmark suites [9] and are increasingly being integrated with external tools to further enhance their functionality [10].
@@
Despite their success in processing and generating human-like language, these models harbor a fundamental limitation: they operate primarily by learning statistical correlations from data, lacking a deep, grounded understanding of the world.[11]This gap becomes particularly apparent when LLMs are confronted with tasks requiring intuitive physical reasoning [12]. While they can often recite physical laws from their training data, they frequently fail to apply these principles to novel scenarios, a limitation that has been extensively documented [13]. Advanced prompting strategies such as Chain-of-Thought [14] and Tree-of-Thoughts [15] have been proposed to elicit more structured reasoning. However, these methods are essentially techniques for organizing the model’s existing knowledge rather than instilling a genuine, predictive understanding of physical dynamics [16]. This deficiency hinders the deployment of LLMs in high-stakes, real-world applications like robotics and autonomous systems, where the ability to predict the consequences of actions is paramount [17, 18]. The challenge is not merely about recognizing objects, but about understanding their interactions, a problem also explored in audio-visual event analysis [19].
@@
The core of the issue lies in the distinction between correlation and causation. Current models excel at the former, but true physical intelligence requires the latter. This has led to a growing interest in building models that are more physically grounded [20, 21]. Efforts in this direction include creating embodied multimodal models [22], developing new benchmarks for physical and event-based reasoning [23, 24, 25], and exploring novel architectures for efficient video analysis, such as unified networks for temporal filtering [26]. This is part of a broader trend in creating generative interactive environments from visual data [27]. Concurrently, research in other domains has focused on interpreting the physical world through diverse non-visual signals. WiFi-based sensing, for example, has proven to be a versatile modality. Foundational work established robust human activity recognition by developing methods to mitigate co-channel interference [28], a crucial step for any real-world interactive system [29]. This reliability is paramount for embodied agents that must ground language in robotic affordances [17]. Building on this, more complex applications like gesture recognition have emerged, where systems must not only be accurate but also handle unknown, open-set gestures to be practical in the real world [30]. The application space for such sensing has expanded dramatically, a trend seen across many forms of human-AI interaction [31], moving into contactless healthcare with systems that can capture pulmonary function [32]. This extends beyond WiFi, with other RF technologies like RFID-based systems also being explored for identifying fine-grained activities such as writing [5]. Such sensing paradigms enable nuanced tasks that require a form of machine commonsense to interpret [33], including unobtrusive emotion recognition [34]. However, the increasing sensitivity of data collected by such pervasive sensing also raises significant security concerns, leading to research into threats like keystroke eavesdropping from commodity devices [35]. These specialized systems, while powerful, often focus on classification rather than building a general, predictive model of the world. The development of such general models is computationally intensive, driving research into distributed learning schemes like federated learning with experience-driven model migration [36]. This is complemented by a focus on algorithmic efficiency through novel architectures [37] and advanced techniques like multi-objective model compression [38].
@@
To bridge this gap, we turn to the concept of ”world models,” which has shown great promise in reinforcement learning for learning predictive models of an environment [39, 40]. We posit that for an LLM to achieve genuine physical reasoning, it must be endowed with an internal, causal world model. This paper introduces a novel framework, Causal World Model Induction (CWMI), designed to explicitly embed a model of causal physics within an LLM. Our approach moves beyond simple next-token prediction by integrating a specialized Causal Physics Module (CPM) and a new training objective that forces the model to learn cause-and-effect relationships from multimodal data. This is inspired by work on relational inductive biases [41] and the need for causal perspectives in machine learning [42, 43]. By training the model to predict the outcomes of hypothetical interventions, we aim to build a more robust and generalizable understanding of physical dynamics, a challenge analogous to suppressing label noise in facial recognition to find the true underlying patterns [44]. Our goal is to create a system that doesn’t just process language, but reasons about the world it describes.
The main contributions of this work are as follows:
@@ II-A Reasoning in Large Language Models
The capacity for reasoning in LLMs has been a subject of intense investigation. Early work focused on the emergent properties of scaled-up models [2]. More recently, research has shifted towards explicit techniques to guide the reasoning process. Chain-of-Thought (CoT) prompting [14] and its derivatives, such as Tree of Thoughts [15], have shown that instructing models to ”think step-by-step” can significantly improve performance on arithmetic, commonsense, and symbolic reasoning tasks. Other approaches aim to augment LLMs with external tools or code execution engines, allowing them to offload complex calculations or verify information, as seen in Toolformer [10] and ViperGPT [45]. Frameworks have been proposed to structure the entire comprehension-reasoning-planning process [46]. However, these methods primarily enhance the application of knowledge already stored in the model’s parameters. They do not fundamentally alter the model’s internal representation of the world, which often lacks the physical grounding necessary for robust, real-world interaction [20]. This is a critical distinction, as a model can articulate a flawless logical chain yet still base it on a physically implausible premise.
@@ II-B World Models and Physical Prediction
The concept of a ”world model” [47], an internal, predictive model of an environment, originated largely in the context of reinforcement learning and robotics [39]. Models like DreamerV3 [40] learn directly from sensory inputs to create a latent space where planning can occur, enabling agents to master diverse and complex tasks from rich data. This includes understanding and predicting fine-grained micro-actions, a challenge addressed by dedicated benchmarking efforts [48]. In parallel, the computer vision community has developed models for explicit physical prediction from visual data. Physics-informed neural networks (PINNs) embed differential equations directly into the learning process [49]. Other work has focused on creating challenging benchmarks like CLEVRER [23] and Physion [25] to evaluate models on their ability to predict the outcomes of physical events. Recently, generative models like GENIE [27] have shown the ability to create entire interactive environments from a single image, demonstrating a powerful form of implicit world modeling. Our work draws inspiration from these approaches but differs in a key aspect: instead of learning from low-level sensory data or within a constrained simulation, we aim to induce a general, causal world model within a pre-trained LLM using a combination of textual descriptions and video observations. This leverages the vast knowledge already present in the LLM while grounding it in causal physical principles. The development of such complex models relies on advances in underlying neural architectures, such as Mamba [37], and efficient learning schemes across heterogeneous networks [36].
@@ II-C Causality and Multimodal Grounding
A central thesis of our work is that true physical reasoning requires a causal understanding. The field of causal inference provides a formal language for this, but applying it to high-dimensional data like text and video is a significant challenge [42]. Some work has explored using counterfactuals for data augmentation [50] or learning causal relationships in a compositional manner [43]. In the multimodal domain, significant effort has been directed at grounding language in perception. This includes pre-training on large-scale video-language datasets [51, 52], developing neuro-symbolic architectures that combine perception with explicit reasoning [53], and creating embodied agents that learn from interaction [54]. The challenge of grounding is also present in other sensing modalities, where researchers aim to prevent attacks on physical-layer authentication [55]. This focus on security and reliability is paramount for embodied agents that must function in noisy, real-world conditions [17]. In a similar vein, others seek to build robust activity recognition systems by developing anti-interference techniques based on phase components [56]. The ability to extract a clean signal from noise is a general problem, with solutions often involving novel neural module communication strategies [57]. The successful application of these robust techniques enables more advanced, target-oriented sensing in cluttered environments for applications like healthcare [58]. Interpreting these signals to understand high-level human states, such as emotions, requires a form of intuitive physics or theory of mind [59]. This is exemplified by systems that fuse WiFi and vision for unobtrusive emotion recognition [34]. Our work directly tackles this by using causal interventions as a supervisory signal, forcing the model to learn not just *what* happens, but *why* it happens. This is analogous to efforts in other fields to learn from heterogeneous data sources to a more reliable representation [60] or perform commonsense reasoning over knowledge graphs [33].
@@ III-A2 Zero-Shot Evaluation
To assess the generalization capability of our model, we perform zero-shot evaluation on the well-established PIQA (Physical Interaction Quality Assessment) benchmark [24]. PIQA is a multiple-choice question-answering dataset that requires commonsense reasoning about how to perform everyday tasks involving physical interaction. It contains 16,113 training examples and 1,838 for the public test set. Crucially, we do not use any of the PIQA training data during the training or fine-tuning of our model to ensure a true zero-shot evaluation setting.
@@ IV Methodology
The central hypothesis of our work is that genuine physical reasoning in artificial intelligence cannot emerge from statistical pattern matching alone. Instead, it requires an internal, predictive, and causal model of the world. Current Large Language Models (LLMs), despite their linguistic prowess [2, 4], fundamentally lack this capability, operating as sophisticated stochastic parrots that can describe but not truly understand physical dynamics [13]. To address this critical gap, we introduce the Causal World Model Induction (CWMI) framework. Our methodology is designed not to simply fine-tune an LLM on physics-related text, but to fundamentally augment it with a new computational module explicitly engineered to simulate cause and effect.
@@ IV-B1 State Representation
A critical design choice is how to represent a physical scene within the CPM’s latent space. Inspired by relational inductive biases in graph networks [41] and object-centric learning, our state vector $S\in\mathbb{R}^{d_{cpm}}$ is structured to implicitly represent the key entities and their properties. While not explicitly factored into distinct object slots in the current implementation, the dimensionality and training objective encourage the model to learn a disentangled representation where different dimensions correspond to properties like position, velocity, mass, and material of the primary objects in the scene. For a scene with up to $N$ objects, each with $P$ properties, the state dimension $d_{cpm}$ would be conceptually related to $N\times P$. The LLM-to-CPM projection layer is tasked with the crucial role of extracting this structured information from the LLM’s dense, holistic embedding.
@@ IV-C2 Causal Intervention Loss ($L_{causal}$)
Physical Meaning: This loss function has a profound physical and causal interpretation. It does not reward the model for correctly predicting the absolute final states $\hat{S}_{final}$ or $\hat{S}^{\prime}_{final}$—that is the job of $L_{pred}$. Instead, $L_{causal}$ focuses exclusively on whether the model correctly predicts the *effect* of the intervention. For example, if changing a ball from rubber to steel (the intervention) causes its final velocity to increase by 5 m/s (the effect), this loss function pushes the model to learn precisely that relationship. It forces the model to isolate the causal link between a specific change in an initial condition and the resulting change in the outcome. By training on a diverse set of such interventions (changing mass, friction, initial velocity, etc.), the CPM learns a robust model of physics that is less susceptible to spurious correlations. This approach is a practical implementation of the ideas from structural causal models [42] and counterfactual reasoning [50] tailored for deep learning models.
@@ IV-D Zero-Shot Inference for Physical QA
Consider a multiple-choice question from a benchmark like PIQA [24]. The input consists of a premise (e.g., ”To dry a shirt, you hang it on a clothesline.”) and two possible solutions (e.g., Sol 1: ”The sun and wind evaporate the water.”; Sol 2: ”The shirt absorbs sunlight, making it heavier.”).
@@ V-A1 Performance on the PIQA Benchmark
We first assessed our model on the Physical Interaction Quality Assessment (PIQA) benchmark [24], which requires commonsense understanding of everyday physical situations. The results, summarized in Table I, demonstrate a clear and significant advantage for our approach. Our CWMI model achieves a zero-shot accuracy of 89.4%, outperforming the base Llama 3 8B model by a substantial margin of 15.2 percentage points. More notably, it also surpasses the reported performance of much larger, proprietary models like GPT-4, which, despite their vast parameter counts and training data, still struggle with the nuances of physical common sense [12, 3].
@@
The performance gap, visualized in the bar chart in Fig. LABEL:fig:barchart_piqa, is telling. Standard LLMs approach physical reasoning as a language modeling problem; they select the answer that is most statistically probable given their training corpus [16]. This makes them vulnerable to ”fluent fallacies”—answers that are linguistically plausible but physically nonsensical. Our CWMI framework, by contrast, approaches the problem as a simulation task. As described in our methodology (Section IV), it uses its Causal Physics Module (CPM) to internally simulate the outcome of each potential solution and selects the one that is most physically coherent. This ability to perform latent-space simulation provides a powerful inductive bias that is absent in standard architectures. While advanced prompting techniques like Chain-of-Thought [14] can elicit better reasoning from base models, they are fundamentally post-hoc strategies that attempt to structure the model’s existing, flawed knowledge. Our work, in contrast, directly addresses the flaw in the underlying knowledge itself by embedding a more robust, causal model of the world. This result validates that for tasks grounded in the physical world, a specialized, model-based reasoning component is superior to relying on the emergent, and often brittle, capabilities of a pure language model.

### v4

@@ I Introduction
The last few years have witnessed a paradigm shift in artificial intelligence, driven by the advent of Large Language Models (LLMs). Foundational models like BERT [Devlin2019BERT] and the GPT series [Brown2020GPT3, Bubeck2023Sparks] have demonstrated remarkable, often emergent, capabilities across a vast spectrum of natural language tasks. Their proficiency is not limited to text; recent multimodal models such as Gemini [Team2024Gemini] and Flamingo [Alayrac2022Flamingo] have extended this power to encompass vision and other modalities, achieving impressive results in areas like visual dialog and video question answering [Zhang2022Survey]. The capabilities of these models are so extensive that they are often evaluated on large, complex benchmark suites [BIG-Bench-Authors2023] and are increasingly being integrated with external tools to further enhance their functionality [Schick2023Toolformer].
@@
Despite their success in processing and generating human-like language, these models harbor a fundamental limitation: they operate primarily by learning statistical correlations from data, lacking a deep, grounded understanding of the world.This gap becomes particularly apparent when LLMs are confronted with tasks requiring intuitive physical reasoning [Li2023GPT4Physics]. While they can often recite physical laws from their training data, they frequently fail to apply these principles to novel scenarios, a limitation that has been extensively documented [valmeekam2022can]. Advanced prompting strategies such as Chain-of-Thought [Wei2022CoT] and Tree-of-Thoughts [Yao2023Tree] have been proposed to elicit more structured reasoning. However, these methods are essentially techniques for organizing the model’s existing knowledge rather than instilling a genuine, predictive understanding of physical dynamics [Shanahan2022Symbolic]. This deficiency hinders the deployment of LLMs in high-stakes, real-world applications like robotics and autonomous systems, where the ability to predict the consequences of actions is paramount [Ahn2022SayCan, Huang2022InnerMonologue]. The challenge is not merely about recognizing objects, but about understanding their interactions, a problem also explored in audio-visual event analysis .
@@
The core of the issue lies in the distinction between correlation and causation. Current models excel at the former, but true physical intelligence requires the latter. This has led to a growing interest in building models that are more physically grounded [Ding2023LLMReasoning, Li2023Symbolic]. Efforts in this direction include creating embodied multimodal models [driess2023palme], developing new benchmarks for physical and event-based reasoning [Yi2020NSDR, Zellers2019PIQA, Bear2021Physion], and exploring novel architectures for efficient video analysis, such as unified networks for temporal filtering . This is part of a broader trend in creating generative interactive environments from visual data [Team2024Genie]. Concurrently, research in other domains has focused on interpreting the physical world through diverse non-visual signals. WiFi-based sensing, for example, has proven to be a versatile modality. Foundational work established robust human activity recognition by developing methods to mitigate co-channel interference , a crucial step for any real-world interactive system [Park2023Agents]. This reliability is paramount for embodied agents that must ground language in robotic affordances [Ahn2022SayCan]. Building on this, more complex applications like gesture recognition have emerged, where systems must not only be accurate but also handle unknown, open-set gestures to be practical in the real world . The application space for such sensing has expanded dramatically, a trend seen across many forms of human-AI interaction [Wu2023MINT], moving into contactless healthcare with systems that can capture pulmonary function . This extends beyond WiFi, with other RF technologies like RFID-based systems also being explored for identifying fine-grained activities such as writing . Such sensing paradigms enable nuanced tasks that require a form of machine commonsense to interpret [Sap2019Atomic], including unobtrusive emotion recognition . However, the increasing sensitivity of data collected by such pervasive sensing also raises significant security concerns, leading to research into threats like keystroke eavesdropping from commodity devices . These specialized systems, while powerful, often focus on classification rather than building a general, predictive model of the world. The development of such general models is computationally intensive, driving research into distributed learning schemes like federated learning with experience-driven model migration . This is complemented by a focus on algorithmic efficiency through novel architectures [Gu2023Mamba] and advanced techniques like multi-objective model compression .
@@
To bridge this gap, we turn to the concept of ”world models,” which has shown great promise in reinforcement learning for learning predictive models of an environment [Ha2018WorldModels, Hafner2023DreamerV3]. We posit that for an LLM to achieve genuine physical reasoning, it must be endowed with an internal, causal world model. This paper introduces a novel framework, Causal World Model Induction (CWMI), designed to explicitly embed a model of causal physics within an LLM. Our approach moves beyond simple next-token prediction by integrating a specialized Causal Physics Module (CPM) and a new training objective that forces the model to learn cause-and-effect relationships from multimodal data. This is inspired by work on relational inductive biases [Battaglia2018GNNs] and the need for causal perspectives in machine learning [Kiciman2024Causal, Dasgupta2021Grounded]. By training the model to predict the outcomes of hypothetical interventions, we aim to build a more robust and generalizable understanding of physical dynamics, a challenge analogous to suppressing label noise in facial recognition to find the true underlying patterns . Our goal is to create a system that doesn’t just process language, but reasons about the world it describes.
@@ II-A Reasoning in Large Language Models
The capacity for reasoning in LLMs has been a subject of intense investigation. Early work focused on the emergent properties of scaled-up models [Brown2020GPT3]. More recently, research has shifted towards explicit techniques to guide the reasoning process. Chain-of-Thought (CoT) prompting [Wei2022CoT] and its derivatives, such as Tree of Thoughts [Yao2023Tree], have shown that instructing models to ”think step-by-step” can significantly improve performance on arithmetic, commonsense, and symbolic reasoning tasks. Other approaches aim to augment LLMs with external tools or code execution engines, allowing them to offload complex calculations or verify information, as seen in Toolformer [Schick2023Toolformer] and ViperGPT [Li2023ViperGPT]. Frameworks have been proposed to structure the entire comprehension-reasoning-planning process [Zhang2023Igniting]. However, these methods primarily enhance the application of knowledge already stored in the model’s parameters. They do not fundamentally alter the model’s internal representation of the world, which often lacks the physical grounding necessary for robust, real-world interaction [Ding2023LLMReasoning]. This is a critical distinction, as a model can articulate a flawless logical chain yet still base it on a physically implausible premise.
@@ II-B World Models and Physical Prediction
The concept of a ”world model” [Lecun2022Path], an internal, predictive model of an environment, originated largely in the context of reinforcement learning and robotics [Ha2018WorldModels]. Models like DreamerV3 [Hafner2023DreamerV3] learn directly from sensory inputs to create a latent space where planning can occur, enabling agents to master diverse and complex tasks from rich data. This includes understanding and predicting fine-grained micro-actions, a challenge addressed by dedicated benchmarking efforts . In parallel, the computer vision community has developed models for explicit physical prediction from visual data. Physics-informed neural networks (PINNs) embed differential equations directly into the learning process [Raissi2019PINNs]. Other work has focused on creating challenging benchmarks like CLEVRER [Yi2020NSDR] and Physion [Bear2021Physion] to evaluate models on their ability to predict the outcomes of physical events. Recently, generative models like GENIE [Team2024Genie] have shown the ability to create entire interactive environments from a single image, demonstrating a powerful form of implicit world modeling. Our work draws inspiration from these approaches but differs in a key aspect: instead of learning from low-level sensory data or within a constrained simulation, we aim to induce a general, causal world model within a pre-trained LLM using a combination of textual descriptions and video observations. This leverages the vast knowledge already present in the LLM while grounding it in causal physical principles. The development of such complex models relies on advances in underlying neural architectures, such as Mamba [Gu2023Mamba], and efficient learning schemes across heterogeneous networks .
@@ II-C Causality and Multimodal Grounding
A central thesis of our work is that true physical reasoning requires a causal understanding. The field of causal inference provides a formal language for this, but applying it to high-dimensional data like text and video is a significant challenge [Kiciman2024Causal]. Some work has explored using counterfactuals for data augmentation [Goyal2019Counterfactuals] or learning causal relationships in a compositional manner [Dasgupta2021Grounded]. In the multimodal domain, significant effort has been directed at grounding language in perception. This includes pre-training on large-scale video-language datasets [Jin2024VILA, Girdhar2023LAVILA], developing neuro-symbolic architectures that combine perception with explicit reasoning [Mao2019NSCL], and creating embodied agents that learn from interaction [Chen2023VIMA]. The challenge of grounding is also present in other sensing modalities, where researchers aim to prevent attacks on physical-layer authentication . This focus on security and reliability is paramount for embodied agents that must function in noisy, real-world conditions [Ahn2022SayCan]. In a similar vein, others seek to build robust activity recognition systems by developing anti-interference techniques based on phase components . The ability to extract a clean signal from noise is a general problem, with solutions often involving novel neural module communication strategies [Mott2019MIC]. The successful application of these robust techniques enables more advanced, target-oriented sensing in cluttered environments for applications like healthcare . Interpreting these signals to understand high-level human states, such as emotions, requires a form of intuitive physics or theory of mind [Piloto2022Intuitive]. This is exemplified by systems that fuse WiFi and vision for unobtrusive emotion recognition . Our work directly tackles this by using causal interventions as a supervisory signal, forcing the model to learn not just *what* happens, but *why* it happens. This is analogous to efforts in other fields to learn from heterogeneous data sources to a more reliable representation or perform commonsense reasoning over knowledge graphs [Sap2019Atomic].
@@ III-A2 Zero-Shot Evaluation
To assess the generalization capability of our model, we perform zero-shot evaluation on the well-established PIQA (Physical Interaction Quality Assessment) benchmark [Zellers2019PIQA]. PIQA is a multiple-choice question-answering dataset that requires commonsense reasoning about how to perform everyday tasks involving physical interaction. It contains 16,113 training examples and 1,838 for the public test set. Crucially, we do not use any of the PIQA training data during the training or fine-tuning of our model to ensure a true zero-shot evaluation setting.
@@ IV Methodology
The central hypothesis of our work is that genuine physical reasoning in artificial intelligence cannot emerge from statistical pattern matching alone. Instead, it requires an internal, predictive, and causal model of the world. Current Large Language Models (LLMs), despite their linguistic prowess [Brown2020GPT3, Team2024Gemini], fundamentally lack this capability, operating as sophisticated stochastic parrots that can describe but not truly understand physical dynamics [valmeekam2022can]. To address this critical gap, we introduce the Causal World Model Induction (CWMI) framework. Our methodology is designed not to simply fine-tune an LLM on physics-related text, but to fundamentally augment it with a new computational module explicitly engineered to simulate cause and effect.
@@ IV-B1 State Representation
A critical design choice is how to represent a physical scene within the CPM’s latent space. Inspired by relational inductive biases in graph networks [Battaglia2018GNNs] and object-centric learning, our state vector $S\in\mathbb{R}^{d_{cpm}}$ is structured to implicitly represent the key entities and their properties. While not explicitly factored into distinct object slots in the current implementation, the dimensionality and training objective encourage the model to learn a disentangled representation where different dimensions correspond to properties like position, velocity, mass, and material of the primary objects in the scene. For a scene with up to $N$ objects, each with $P$ properties, the state dimension $d_{cpm}$ would be conceptually related to $N\times P$. The LLM-to-CPM projection layer is tasked with the crucial role of extracting this structured information from the LLM’s dense, holistic embedding.
@@ IV-C2 Causal Intervention Loss ($L_{causal}$)
Physical Meaning: This loss function has a profound physical and causal interpretation. It does not reward the model for correctly predicting the absolute final states $\hat{S}_{final}$ or $\hat{S}^{\prime}_{final}$—that is the job of $L_{pred}$. Instead, $L_{causal}$ focuses exclusively on whether the model correctly predicts the *effect* of the intervention. For example, if changing a ball from rubber to steel (the intervention) causes its final velocity to increase by 5 m/s (the effect), this loss function pushes the model to learn precisely that relationship. It forces the model to isolate the causal link between a specific change in an initial condition and the resulting change in the outcome. By training on a diverse set of such interventions (changing mass, friction, initial velocity, etc.), the CPM learns a robust model of physics that is less susceptible to spurious correlations. This approach is a practical implementation of the ideas from structural causal models [Kiciman2024Causal] and counterfactual reasoning [Goyal2019Counterfactuals] tailored for deep learning models.
@@ IV-D Zero-Shot Inference for Physical QA
Consider a multiple-choice question from a benchmark like PIQA [Zellers2019PIQA]. The input consists of a premise (e.g., ”To dry a shirt, you hang it on a clothesline.”) and two possible solutions (e.g., Sol 1: ”The sun and wind evaporate the water.”; Sol 2: ”The shirt absorbs sunlight, making it heavier.”).
@@ V-A1 Performance on the PIQA Benchmark
We first assessed our model on the Physical Interaction Quality Assessment (PIQA) benchmark [Zellers2019PIQA], which requires commonsense understanding of everyday physical situations. The results, summarized in Table I, demonstrate a clear and significant advantage for our approach. Our CWMI model achieves a zero-shot accuracy of 89.4%, outperforming the base Llama 3 8B model by a substantial margin of 15.2 percentage points. More notably, it also surpasses the reported performance of much larger, proprietary models like GPT-4, which, despite their vast parameter counts and training data, still struggle with the nuances of physical common sense [Li2023GPT4Physics, Bubeck2023Sparks].
@@
The performance gap, visualized in the bar chart , is telling. Standard LLMs approach physical reasoning as a language modeling problem; they select the answer that is most statistically probable given their training corpus [Shanahan2022Symbolic]. This makes them vulnerable to ”fluent fallacies”—answers that are linguistically plausible but physically nonsensical. Our CWMI framework, by contrast, approaches the problem as a simulation task. As described in our methodology (Section IV), it uses its Causal Physics Module (CPM) to internally simulate the outcome of each potential solution and selects the one that is most physically coherent. This ability to perform latent-space simulation provides a powerful inductive bias that is absent in standard architectures. While advanced prompting techniques like Chain-of-Thought [Wei2022CoT] can elicit better reasoning from base models, they are fundamentally post-hoc strategies that attempt to structure the model’s existing, flawed knowledge. Our work, in contrast, directly addresses the flaw in the underlying knowledge itself by embedding a more robust, causal model of the world. This result validates that for tasks grounded in the physical world, a specialized, model-based reasoning component is superior to relying on the emergent, and often brittle, capabilities of a pure language model.